# Solvable Tuple Patterns and Their Applications to Program Verification

### v1

@@ A.1.7 Alternative Proof of Minimality
Proof of Lemma A.5 (for the extension with postfix).
We need to consider only the cases where $t\leadsto t_{1}$ uses PR-Prefix or PR-CPrefix, and $t\leadsto t_{2}$ uses PR-Postfix or PR-CPostfix. Below we discuss only the case where $t\leadsto t_{1}$ uses PR-Prefix and $t\leadsto t_{2}$ uses PR-Postfix; the cases where PR-CPrefix or PR-CPostfix is used are similar.
[…]
  • Case where $t=(p_{2}p_{1}^{\prime},p_{2},\ldots,p_{n})\leadsto(p_{1}^{\prime},p_{2},\ldots,p_{n})=t_{1}$ and $t=(p_{1}^{\prime\prime}p_{3},p_{2},p_{3},\ldots,p_{n})\leadsto(p_{1}^{\prime\prime},p_{2},p_{3},\ldots,p_{n})=t_{2}$. In this case, $p_{2}p_{1}^{\prime}=p_{1}^{\prime\prime}p_{3}$; so, either (i) $p_{1}^{\prime\prime}=p_{2}p_{1}^{\prime\prime\prime}$ and $p_{1}^{\prime}=p_{1}^{\prime\prime\prime}p_{3}$, or (ii) $p_{2}=p_{1}^{\prime\prime}p_{2}^{\prime}$ and $p_{3}=p_{2}^{\prime}p_{1}^{\prime}$. In case (i), we have:
    In case (ii), we have:
  • Case where $t=(p_{2}p_{1}^{\prime},p_{2},\ldots,p_{n})\leadsto(p_{1}^{\prime},p_{2},\ldots,p_{n})=t_{1}$ and $t=(p_{1},p_{2}^{\prime}p_{3},p_{3},\ldots,p_{n})\leadsto(p_{1},p^{\prime}_{2},p_{3},\ldots,p_{n})=t_{2}$. In this case, we have $p_{1}=p_{2}p_{1}^{\prime}$ and $p_{2}=p_{2}^{\prime}p_{3}$. Then, we have:
    as required.
A.2.2 The extension in Section 2.3.2.
We informally show that all the main properties are preserved by the extension above. In fact, the rules R-RPrefix and R-RPostfix introduced above can be emulated by R-Postfix and R-Prefix, if we augment the data $M$ by adding, for each column M[][i]M[*][i] of $M$, a new column consisting of the reverse of elements of M[][i]M[*][i]. For example, we have:
@@
$\longrightarrow((x_{1},y_{1}x^{\prime}_{2},y_{1},y^{\prime}_{2}x_{1}),[\left(\begin{array}[]{rrrr}ab&\ c&\ ba&\ c\\ \end{array}\right)/(x_{1},x^{\prime}_{2},y_{1},y^{\prime}_{2})])$ (R-Postfix)

### v2

@@ A.1.3 Correspondence between $\longrightarrow$ and $\leadsto$
$\rho\models(w_{1},\ldots,w_{n})\in\mathcal{L}(t)\mbox{ if }(\rho w_{1},\ldots,\rho w_{n})\in\mathcal{L}(t)$
$\rho\models(w_{1},\ldots,w_{n})\not\in\mathcal{L}(t)\mbox{ if }(\rho w_{1},\ldots,\rho w_{n})\not\in\mathcal{L}(t)$
$\rho\models\varphi_{1}\land\varphi_{2}\mbox{ if }\rho\models\varphi_{1}\mbox{ and }\rho\models\varphi_{2}$
$\rho\models\varphi_{1}\lor\varphi_{2}\mbox{ if }\rho\models\varphi_{1}\mbox{ or }\rho\models\varphi_{2}.$
We say a formula $\varphi$ is satisfiable if $\rho\models\varphi$ for some $\rho$. Without subformulas $\widetilde{w}\in\mathcal{L}(t)$ and $\widetilde{w}\not\in\mathcal{L}(t)$, it is known that the satisfiability of quantifier-free formulas is decidable [41, 51, 32].
Proof of Theorem 3.10.
By the decidability of the satisfiability of word equations [41, 51, 32], it suffices to show that atomic formulas of the form $w_{1}\neq w_{2}$, $(w_{1},\ldots,w_{k})\in\mathcal{L}(t)$, or $(w_{1},\ldots,w_{k})\not\in\mathcal{L}(t)$ can be encoded into formulas using only word equations. We can assume, without loss of generality, that the alphabet $\Sigma$ is finite: note that if $\varphi$ is satisfiable for an infinite alphabet, then $\varphi$ is also satisfiable for $\Sigma_{\varphi}\cup\{a,b\}$ where $\Sigma_{\varphi}$ is the set of letters occurring in $\varphi$ and $a,b$ are distinct letters not occurring in $\varphi$, and vice versa. First, we encode the relation $w_{1}\not\preceq w_{2}$, which means $w_{1}$ is not a prefix of $w_{2}$.” It can be expressed by $(\bigvee_{a,b\in\Sigma,a\neq b}w_{1}=xay\land w_{2}=xbz)\lor\bigvee_{a\in\Sigma}w_{1}=w_{2}ax$, where $x,y,z$ are fresh variables. (This encoding of $\not\preceq$ is standard; see, e.g. [14, 32].) We can then express $w_{1}\neq w_{2}$ as $w_{1}\not\preceq w_{2}\lor w_{2}\not\preceq w_{1}$.
The encoding of $(w_{1},\ldots,w_{k})\in\mathcal{L}(t)$ is straightforward. Suppose $t=(p_{1},\ldots,p_{\ell})$. If $k\neq\ell$, then $(w_{1},\ldots,w_{k})\in\mathcal{L}(t)$ can be replaced with $\mathit{false}{}$ (which can be expressed as $a=\epsilon$). If $k=\ell$, then $(w_{1},\ldots,w_{k})\in\mathcal{L}(t)$ is expressed as $w_{1}=p^{\prime}_{1}\land\cdots\land w_{k}=p^{\prime}_{k}$, where $(p^{\prime}_{1},\ldots,p^{\prime}_{k})$ are obtained from $t$ by $\alpha$-renaming, so that the variables in $p^{\prime}_{1},\ldots,p^{\prime}_{k}$ are fresh.
Now we consider $(w_{1},\ldots,w_{k})\not\in\mathcal{L}(t)$, where $t=(p_{1},\ldots,p_{\ell})$. If $k\neq\ell$, then it can be replaced with $\mathit{true}{}$ (which can be expressed as $\epsilon=\epsilon$). Thus, we assume $k=\ell$. Since $t$ is solvable, there exists a reduction sequence $t\leadsto^{*}(x_{1},\ldots,x_{m})$. We encode the formula by induction on the length $n$ of the reduction sequence. If $n=0$, then $\mathcal{L}(t)=\mathcal{L}(x_{1},\ldots,x_{k})=\Sigma^{*}\times\cdots\times\Sigma^{*}$. Thus, $(w_{1},\ldots,w_{k})\not\in\mathcal{L}(t)$ can be replaced with $\mathit{false}$. If $n>0$, then $t\leadsto t^{\prime}\leadsto^{*}(x_{1},\ldots,x_{m})$. We perform case analysis on the rule used for deriving $t\leadsto t^{\prime}$. Since the other cases are similar, we discuss only the cases for PR-Prefix and PR-Epsilon.
  • Case PR-Prefix: In this case, there exist $i,j$ such that $t^{\prime}=(p^{\prime}_{1},\ldots,p^{\prime}_{k})$ with $p_{j}=p_{i}p^{\prime}_{j}$ and $p^{\prime}_{\ell}=p_{\ell}$ for all $\ell\neq j$. Thus, $(w_{1},\ldots,w_{k})\not\in\mathcal{L}(t)$ can be replaced with $w_{i}\not\preceq w_{j}\lor(w_{j}=w_{i}x\land(w^{\prime}_{1},\ldots,w^{\prime}_{k})\not\in\mathcal{L}(t^{\prime}))$, where $x$ is a fresh variable, $w^{\prime}_{j}=x$, and $w^{\prime}_{\ell}=w_{\ell}$ for all $\ell\neq j$. By the induction hypothesis, $(w^{\prime}_{1},\ldots,w^{\prime}_{k})\not\in\mathcal{L}(t^{\prime})$ can also be encoded into word equations.
  • Case PR-Epsilon: In this case, there exists $j$ such that $p_{j}=\epsilon$ and $t^{\prime}=(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{k})$. Thus, $(w_{1},\ldots,w_{k})\not\in\mathcal{L}(t)$ can be replaced with $w_{j}\neq\epsilon\lor(w_{1},\ldots,w_{j-1},w_{j+1},\ldots,w_{k})\not\in\mathcal{L}(t^{\prime})$. By the induction hypothesis, $(w^{\prime}_{1},\ldots,w^{\prime}_{k})\not\in\mathcal{L}(t^{\prime})$ can also be encoded into word equations.
@@
A.2 Additional Definitions and Proofs for Section 4
A.2.1 The Extension in Section 4.1
First, we extend $\sim$ (used in Lemma A.6) to the least equivalence relation closed under not only the permutations and but also the rule $(p_{1},\ldots,p_{k})\sim(p_{1},\ldots,p_{i-1},p_{i}^{R},p_{i+1},\ldots,p_{k})$, and slightly relax the requirement for solvability: a tuple pattern $t$ is solvable if $t\leadsto^{*}\sim(x_{1},\ldots,x_{n})$ for some mutually distinct variables $x_{1},\ldots,x_{n}$. Thus, $(x^{R},y)$ is also solvable, as $(x^{R},y)\sim(x,y)$. Without the reverse pattern constructor, the definition of solvability is the same.
All the main properties are preserved by the extension of patterns with the reverse constructor. This can be intuitively understood by noting that the rules R-RPrefix and R-RSuffix can be emulated by R-Prefix and R-Suffix, if we augment the data $M$ by adding, for each column M[][i]M[*][i] of $M$, a new column consisting of the reverse of elements of M[][i]M[*][i]. For example,
@@ A.2 Additional Definitions and Proofs for Section 4
$\longrightarrow((x_{1},y_{1}x^{\prime}_{2},y_{1},y^{\prime}_{2}x_{1}),[\left(\begin{array}[]{rrrr}ab&\ c&\ ba&\ c\\ \end{array}\right)/(x_{1},x^{\prime}_{2},y_{1},y^{\prime}_{2})])$ (R-Suffix)